OpenStudy (anonymous):

A piano recital is scheduled to be ninety minutes long. The piano teacher will play for ten minutes. During the remaining time, the twenty piano students will get equal time for their performance. How long will each piano student perform? Let m represent the number of minutes each piano student performs. ninety plus twenty times m equals ten twenty times m plus ten equals ninety ten minus twenty times m equals ninety twenty times m equals one hundred

OpenStudy (anonymous):

each performer gets 4 minutes because if there is 90 minutes total - 10 is 80 and 80/20 is 4

OpenStudy (anonymous):

ok

OpenStudy (mathmale):

First of all, this problem involves a linear function: y = m*x + b. You are given y, b and x: y=90 minutes; b=10 minutes; and x=20. Find the slope, m. How?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then wat

OpenStudy (mathmale):

I've already suggested that you solve the above equation for the slope, m. Try it, please. Show me what you've done. Then I'd be happy to give you feedback on your work.

OpenStudy (mathmale):

First of all, this problem involves a linear function: y = m*x + b. You are given y, b and x: y=90 minutes; b=10 minutes; and x=20. Find the slope, m. How?